# Theory of dynamical superradiance in organic materials

### v1

@@ Abstract
Keywords: vibrational coupling; strong light-matter coupling; collective light emission; cavity quantum electrodyamics; open quantum systems; master equations
@@ I Introduction
If the emitters coupled to the cavity mode are treated as ideal two-level systems, and one makes the rotating wave approximation, then the combined light-matter system can be described with the Tavis–Cummings (TC) model [33]. In this case, the behavior of dynamical superradiance is well-studied. A mean-field solution of this model predicts a train of hyperbolic secant pulses for the number of photons in the cavity starting from a fully inverted initial state [20], and finite-size quantum corrections to this solution have been considered in Refs. [34, 35]. The same dynamics have also been studied in the context of cold gases [36, 37, 38, 39], semiconductor microcavities [24, 40], impurity spins coupled to a microwave resonator [41], and free-electron lasers [42]. There further exist studies of complementary initial states, where all emitters are in the ground state and the photon mode is in a Fock state  [43] or a coherent state [44]. Studies of the dynamics in low-excitation subspaces are also given in Refs. [45, 46].
Approximating physical emitters as ideal two-level systems is not typically realistic; furthermore, additional degrees of freedom may lead to richer light-matter coupling phenomena. For example, emitters in solid-state environments are subject to coupling to phonon modes and other sources of dephasing. With organic molecules in particular, there are non-trivial effects of vibrational coupling to the electronic degrees of freedom [47, 48, 49], unless one explicitly works to suppress vibrational excitations [50]. In this work, we extend the treatment of dynamical SR from two-level systems to organic molecules, including vibrational coupling. Organic emitters have been used in experiments demonstrating strong light-matter coupling, polariton lasing, and Bose-Einstein condensation [51, 52, 53, 54, 55]. Hence, they are a promising platform for harnessing collective effects in light-matter coupling.
We consider two different models to describe the influence of the environment of vibrational modes. In the first, the vibrational modes are treated as a Markovian bath for the emitters, through a local dephasing term in a Lindblad master equation, see Fig. 1(a)i. In the second, we include vibrational modes explicitly in the system Hamiltonian, giving the Holstein–Tavis–Cummings (HTC) model [56, 47], Fig. 1(a)ii. We will see that, while in both descriptions vibronic coupling suppresses the extent and duration of the dynamical SR, there are significant differences in behavior, particularly when the vibronic coupling becomes strong.
@@
Experiments, however, have significantly larger numbers of emitters, and so other methods are required to reach large $N$. For this, we use a standard mean-field approach (assuming factorization of the molecular and cavity degrees of freedom), along with a second-order cumulant approach that captures leading-order corrections to the mean-field theory. Such mean-field and cumulant approaches can generally be expected to match the exact solution in the thermodynamic limit $N\rightarrow\infty$ [65, 66, 67], however, this is not always the case [68, 14, 69, 70, 71]. By comparing the exact solution to mean-field and second-order cumulant solutions, we show that the early-time behavior of dynamical SR in the TC model can be captured with both of these approximate methods if the process starts in an SR initial state ($\theta\neq 0$ in Fig. 1(b)), not in the SF ($\theta=0$) one. Since the second-order cumulant solution converges to the mean-field solution in the large $N$ limit, this motivates using mean-field theory to study dynamical SR with vibrationally dressed molecules in the HTC model. We find that strong vibrational coupling destroys dynamical SR, which is most clearly seen in the suppression of the buildup of electronic coherence. However, for weak and moderate vibrational coupling—closer to the typical values in organic molecules—SR is still possible. Interestingly, the vibrational coupling can even assist superradiance when the cavity frequency is smaller than the electronic transition frequency.
@@ II.2 Numerical method
Code for the numerical scheme described is available in the PIBS (Permutational Invariance Block Solver) Python package. This method allows one to solve Eq. (2) with initial conditions given in Eq. (4) up to $N=140$ emitters. Currently, only the solution for the diagonal block $\nu=\nu^{\prime}$ is implemented in the PIBS code, but an extension to off-diagonal blocks is straightforward.
@@ II.3 Benchmarking Cumulant approaches
In Fig. 2, we compare the cumulant approaches to the exact solution for different numbers of emitters and initial conditions, by plotting the normalised average number of photons in the cavity $\langle\hat{n}\rangle/N=\langle\hat{a}^{\dagger}\hat{a}\rangle/N$ as a function of time. For brevity, we refer to the first-order cumulant approximation as mean field and to the second-order cumulant approximation simply as cumulants. We do not show a mean-field solution for SF initial conditions, as the mean-field solution in that case is just zero for all times if we start from the symmetric initial state. We also note that in the mean-field treatment, $\langle\hat{n}\rangle/N$ becomes independent of $N$.
@@
Figures 2(b) and 2(d) compare the exact solution to cumulants and mean field for an SR initial state with $\theta=\pi/4$. Generally, the cumulant solution matches the exact solution better than the mean-field solution. Moreover, for increasing $N$, the early time period where cumulants match the exact solution increases. This indicates that for SR initial states, symmetry-broken cumulants capture the early-time behavior well.
@@ III.1 Model description
The vibrational mode of the $i^{\text{th}}$ molecule is treated as a harmonic oscillator with annihilation and creation operators $\hat{b}_{i}$ and $\hat{b}_{i}^{\dagger}$. The coupling strength between the vibrations and the electronic degrees of freedom is set by the Huang-Rhys parameter $S$. Including the same dissipation processes as before (see Eq. (3111We note that since the dominant vibrational mode is included in the system description in the HTC model, the dephasing terms in Eq. (10) and Eq. (2) have different meanings in principle. However, since we only set $\gamma_{\phi}>0$ in case $S=0$, the two dephasing rates are equivalent.), and adding thermalization processes for the vibrational modes, the master equation is
@@ III.2 Results
To characterize the extent to which the vibrational modes destroy superradiance, we additionally calculate the electronic coherence $|\langle\hat{\sigma}^{+}\rangle|\equiv|\langle\hat{\sigma}_{i}^{+}\rangle|$ for the same sets of parameters used in Figs. 4(a) and 4(d), see Figs. 4(b) and 4(e). Since we use SR initial conditions, the coherence at $t=0$ is non-zero and given by $|\langle\hat{\sigma}^{+}\rangle|=\sin(\theta)/2$. Figure 4(b) shows that the coherence increases exponentially at early times for small $S$, similar to the exponential rise in the photon number in Fig. 4(a). For $S\gtrsim 0.3$, the exponential increase of the coherence is inhibited, and there are regions where the coherence decreases before it has reached its maximum. In Appendix A, we show that increasing the initial amount of coherence (increasing $\theta$) tends to make the initial rise of photon number faster and less dependent on the vibrational coupling.
@@
Such behavior can be related to discussions of how any processes that break spin conservation destroy free-space superradiance [83, 84, 85]. As noted in those works, when the dynamics approaches the equator—specifically when $|\langle\sigma^{z}\rangle|\lesssim 1/\sqrt{N}$—simple counting arguments show there are many more accessible subradiant states (quantum states with $\langle\hat{J}^{2}\rangle\ll N$) than superradiant states. As such, any process that can change the modulus of $\hat{J}$ will lead to transitions to these subradiant states, reducing $\langle\hat{J}^{2}\rangle$ and ultimately destroying the superradiance. Notably, these processes only become relevant as one approaches the equator: when $\langle\sigma^{z}\rangle\simeq 1$, scattering to subradiant states is suppressed if the scattering does not directly change $\langle\sigma^{z}\rangle$. References [83, 84, 85] discuss how direct interactions between different emitters, breaking the permutation symmetry, have such an effect. Here we see similar behavior driven by individual dephasing or vibrational modes.
@@
In experiments, the cavity frequency can typically be readily varied; our calculations predict interesting phenomena when varying the cavity detuning $\Delta=\omega_{\mathrm{c}}-\omega_{0}$, with a clear difference between pure dephasing and having coupling to a vibrational mode. The dynamics of the photon mode in the HTC model is plotted for $\Delta=-0.15\,\mathrm{eV}$ in Fig. 5(a) and for $\Delta=-0.3\,\mathrm{eV}$ in Fig. 5(b). Comparing to the case of zero detuning in Fig. 4(a), we see that a negative detuning favors the emission of photons for larger values of $S$. This is seen most clearly in Fig. 5(b), where the curves up to $S=0.4$ almost collapse into one curve at early times. In the inset of Fig. 5(b), we observe that the photon number rises faster for $0.1\leq S\leq 0.4$ than for $S=0$. This can be explained by first noting that, for $\omega_{\nu}=0.15\,\mathrm{eV}$ and $T=0.026\,\mathrm{eV}$ used here, the phonons are almost fully in the ground state at $t=0$. Further, for a negative detuning, the cavity frequency $\omega_{\mathrm{c}}=\omega_{0}+\Delta$ is smaller than the level splitting of the electronic ground and excited $\omega_{0}$. Thus, the cavity frequency matches transitions from the excited state manifold with zero vibrational excitations to the ground-state manifold with vibrational excitations. As $S$ increases, the most probable number of vibrational excitations created in an optical transition increases. As such, one sees an optimal value of $S$ in this regime where processes exciting one vibrational mode become dominant. In contrast, if the vibrational mode is treated as a pure dephasing term, such a reordering phenomenon does not occur.
@@
The risetime behavior for different values of $\theta$ is discussed in Appendix A. We find that there is an initial rapid rise to a higher photon number for large $\theta$ than for a small one. That is, the photon number rise is faster for a smaller amount of electronic excitation (smaller inversion). This is a curious behavior, opposite, for example, to usual pulsed lasing, where the pulse build-up time becomes shorter for a higher amount of excitation [55]. This is because coherence plays a crucial role in SR, and larger $\theta$ (up to $\pi/2$) means higher initial coherence.
@@ IV Conclusions
Using our exact method, we benchmarked mean-field and second-order cumulant solutions, verifying their accuracy in capturing early-time dynamics when the initial state has non-zero initial coherence. This justified the application of mean-field theory to study the effect of the vibrational environment of organic emitters on the dynamics. Here we compared two models: a Holstein–Tavis–Cummings model including a discrete vibrational mode for each emitter, and a simplified model where the vibronic coupling is replaced by Markovian pure dephasing. We found that for small vibrational coupling or dephasing, both models predict an exponential rise of the photon number at early times. In contrast, strong vibrational coupling leads to qualitatively different photon dynamics in the case of vibrational dressing, which the pure dephasing model cannot capture. Analysis of the electronic coherence revealed that sufficiently strong coupling to vibrations (roughly, $S\geq 0.3$) suppresses dynamical superradiance. As the values of $S$ in typical organic molecules are at a maximum around $0.1$, dynamical superradiance should be possible. On the other hand, the maximal values of the order $0.1$ are large enough to see specific effects of the vibrational coupling, such as the asymmetry of the risetime as a function of cavity detuning.
@@
Data availability: The datasets generated and analysed during the current study are available in the Zenodo repository at https://doi.org/10.5281/zenodo.16994467.
Appendix A Influence of $\theta$ on dynamical SR

### v2

@@ I Introduction
If the emitters coupled to the cavity mode are treated as ideal two-level systems, and one makes the rotating wave approximation, then the combined light-matter system can be described with the Tavis–Cummings (TC) model [33]. In this case, the behavior of dynamical SR is well-studied. A mean-field solution of this model predicts a train of hyperbolic secant pulses for the number of photons in the cavity starting from a fully inverted initial state [20], and finite-size quantum corrections to this solution have been considered in Refs. [34, 35]. The same dynamics have also been studied in the context of cold gases [36, 37, 38, 39], semiconductor microcavities [24, 40], impurity spins coupled to a microwave resonator [41], and free-electron lasers [42]. There further exist studies of complementary initial states, where all emitters are in the ground state and the photon mode is in a Fock state [43] or a coherent state [44]. Studies of the dynamics in low-excitation subspaces are also given in Refs. [45, 46].
Approximating physical emitters as ideal two-level systems is, however, not typically realistic; furthermore, additional degrees of freedom may lead to richer light-matter coupling phenomena. For example, emitters in solid-state environments are subject to coupling to phonon modes and other sources of dephasing. With organic molecules in particular, there are non-trivial effects of vibrational coupling to the electronic degrees of freedom [47, 48, 49], unless one explicitly works to suppress vibrational excitations [50].
Organic emitters have been used in experiments demonstrating strong light-matter coupling, polariton lasing, and Bose-Einstein condensation [51, 52, 53, 54, 55]. Hence, they are a promising platform for harnessing collective effects in light-matter coupling. Motivated by the necessity to model realistic emitters for state-of-the-art experiments, we extend in this work the treatment of dynamical SR from two-level systems [19, 20, 21, 22, 23, 24] to organic molecules, by explicitly incorporating vibrational modes in the system description.
As a first approximation, the effect of a vibrational environment on the electronic degrees of freedom can be treated as a pure dephasing term in a Markovian master equation, see Fig. 1(a)i. This is a phenomenological model aiming to capture the decoherence induced by a vibrational bath that weakly couples to the electronic states, but crucially cannot account for realistic electron-phonon interactions. To go beyond the pure dephasing model, we include one vibrational mode per molecule explicitly in the system Hamiltonian, giving the Holstein–Tavis–Cummings (HTC) model [56, 47, 57, 58, 59, 60], Fig. 1(a)ii. The main advancement of this work is to study dynamical SR in the HTC model. We will see that, while in both the phenomenological model and the HTC model vibrational coupling suppresses the extent and duration of the dynamical SR, there are significant differences in behavior, particularly when the vibronic coupling becomes strong.
@@
Experiments, however, have significantly larger numbers of emitters, and so other methods are required to reach large $N$. For this, we use a standard mean-field approach (assuming factorization of the molecular and cavity degrees of freedom), along with a second-order cumulant approach that captures leading-order corrections to the mean-field theory. Such mean-field and cumulant approaches can generally be expected to match the exact solution in the thermodynamic limit $N\rightarrow\infty$ [69, 70, 71], however, this is not always the case [72, 14, 73, 74, 75]. By comparing the exact solution to mean-field and second-order cumulant solutions, we show that the early-time behavior of dynamical SR in the TC model can be captured with both of these approximate methods if the process starts in an SR initial state ($\theta\neq 0$ in Fig. 1(b)), not in the SF ($\theta=0$) one. Since the second-order cumulant solution converges to the mean-field solution in the large $N$ limit, this motivates using mean-field theory to study dynamical SR with vibrationally dressed molecules in the HTC model.
Cumulant approaches have previously been used to study polariton lasing in the HTC model [76, 77]. However, we note that these works use a reduced set of cumulant equations where terms that vanish due to U(1) symmetry are omitted. Whilst appropriate for studying dynamics from homogeneous initial conditions, such an approach cannot be used to describe SR initial conditions where there is an initial breaking of this symmetry.
We point out another related work [78] where mean-field theory was combined with the process-tensor formalism to solve the HTC model with a continuum of vibrational modes, treated as a non-Markovian bath. In the present work we model the situation where there is one (or a few) vibrational modes strongly coupled to the electronic state, and show how higher-order correlations are included via cumulants.
Our key findings are that strong vibronic coupling destroys dynamical SR, which is most clearly seen in the suppression of the buildup of electronic coherence. However, for weak and moderate vibronic coupling—closer to the typical values in organic molecules—SR is still possible. Interestingly, the vibrational coupling can even assist superradiance when the cavity frequency is smaller than the electronic transition frequency (negative cavity detuning). While a pure dephasing model can account for the suppression of SR, it fails to capture the dependence on the cavity detuning predicted in the HTC model, which is a measurable effect in experiments.
@@ II.2 Numerical method
Code for the numerical scheme described is available in the PIBS (Permutational Invariance Block Solver) Python package [86]. This method allows one to solve Eq. (2) with initial conditions given in Eq. (4) up to $N=140$ emitters. Currently, only the solution for the diagonal block $\nu=\nu^{\prime}$ is implemented in the PIBS code, but an extension to off-diagonal blocks is straightforward.
@@ II.3 Benchmarking Cumulant approaches
In Fig. 2, we compare the cumulant approaches to the exact solution for different numbers of emitters and initial conditions, by plotting the normalised average number of photons in the cavity $\langle\hat{n}\rangle/N=\langle\hat{a}^{\dagger}\hat{a}\rangle/N$ as a function of time. For brevity, we refer to the first-order cumulant approximation as mean field and to the second-order cumulant approximation simply as cumulants. We do not show a mean-field solution for SF initial conditions, as the mean-field solution in that case is just zero for all times if we start from the symmetric initial state. We also note that in the mean-field treatment, $\langle\hat{n}\rangle/N=|\langle\hat{a}\rangle|^{2}/N$ becomes independent of $N$.
@@
Figures 2(b) and 2(d) compare the exact solution to cumulants and mean field for an SR initial state with $\theta=\pi/4$. Generally, the cumulant solution matches the exact solution better than the mean-field solution. Moreover, for increasing $N$, the early time period where cumulants match the exact solution increases. This indicates that for SR initial states, symmetry-broken cumulants capture the early-time behavior well. We provide quantitative evidence that this is the case in Appendix A, where we show that both mean-field and cumulants converge monotonically towards the exact solution for $N$ up to $100$.
@@ III.1 Model description
The vibrational mode of the $i^{\text{th}}$ molecule is treated as a harmonic oscillator with annihilation and creation operators $\hat{b}_{i}$ and $\hat{b}_{i}^{\dagger}$. The coupling strength between the vibrations and the electronic degrees of freedom is set by the Huang-Rhys parameter $S$. Including the same dissipation processes as before (see Eq. (3)  [^1] We note that since the dominant vibrational mode is included in the system description in the HTC model, the dephasing terms in Eq. (10) and Eq. (2) have different meanings in principle. However, since we only set $\gamma_{\phi}>0$ in case $S=0$, the two dephasing rates are equivalent.), and adding thermalization processes for the vibrational modes, the master equation is
@@ III.2 Results
To characterize the extent to which the vibrational modes destroy superradiance, we additionally calculate the electronic coherence $|\langle\hat{\sigma}^{+}\rangle|\equiv|\langle\hat{\sigma}_{i}^{+}\rangle|$ for the same sets of parameters used in Figs. 4(a) and 4(d), see Figs. 4(b) and 4(e). Since we use SR initial conditions, the coherence at $t=0$ is non-zero and given by $|\langle\hat{\sigma}^{+}\rangle|=\sin(\theta)/2$. Figure 4(b) shows that the coherence increases exponentially at early times for small $S$, similar to the exponential rise in the photon number in Fig. 4(a). For $S\gtrsim 0.3$, the exponential increase of the coherence is inhibited, and there are regions where the coherence decreases before it has reached its maximum. In Appendix B, we show that increasing the initial amount of coherence (increasing $\theta$) tends to make the initial rise of photon number faster and less dependent on the vibrational coupling, and in Appendix C we show that the threshold of SR at a certain value of $S$ (here at $S\approx 0.3$) emerges as a competition between the light-matter coupling $g\sqrt{N}$ and the vibronic coupling $\omega_{\nu}\sqrt{S}$.
@@
Such behavior can be related to discussions of how any processes that break spin conservation destroy free-space superradiance [91, 92, 93]. As noted in those works, when the dynamics approaches the equator—specifically when $|\langle\hat{\sigma}^{z}\rangle|\lesssim 1/\sqrt{N}$—simple counting arguments show there are many more accessible subradiant states (quantum states with $\langle\hat{J}^{2}\rangle\ll N$) than superradiant states. As such, any process that can change the modulus of $\hat{J}$ will lead to transitions to these subradiant states, reducing $\langle\hat{J}^{2}\rangle$ and ultimately destroying the superradiance. Notably, these processes only become relevant as one approaches the equator: when $\langle\hat{\sigma}^{z}\rangle\simeq 1$, scattering to subradiant states is suppressed if the scattering does not directly change $\langle\hat{\sigma}^{z}\rangle$. References [91, 92, 93] discuss how direct interactions between different emitters, breaking the permutation symmetry, have such an effect. Here we see similar behavior driven by individual dephasing or vibrational modes.
@@
In experiments, the cavity frequency can typically be readily varied; our calculations predict interesting phenomena when varying the cavity detuning $\Delta=\omega_{\mathrm{c}}-\omega_{0}$, with a clear difference between pure dephasing and having coupling to a vibrational mode. The dynamics of the photon mode in the HTC model is plotted for $\Delta=-0.15\,\mathrm{eV}$ in Fig. 5(a) and for $\Delta=-0.3\,\mathrm{eV}$ in Fig. 5(b). Compared to the case of zero detuning in Fig. 4(a), we see that a negative detuning favors the emission of photons for larger values of $S$. This is seen most clearly in Fig. 5(b), where the curves up to $S=0.4$ almost collapse into one curve at early times. In the inset of Fig. 5(b), we observe that the photon number rises faster for $0.1\leq S\leq 0.4$ than for $S=0$. This can be explained by first noting that, for $\omega_{\nu}=0.15\,\mathrm{eV}$ and $T=0.026\,\mathrm{eV}$ used here, the phonons are almost fully in the ground state at $t=0$. Further, for a negative detuning, the cavity frequency $\omega_{\mathrm{c}}=\omega_{0}+\Delta$ is smaller than the level splitting of the electronic ground and excited $\omega_{0}$. Thus, the cavity frequency matches transitions from the excited state manifold with zero vibrational excitations to the ground-state manifold with vibrational excitations. As $S$ increases, the most probable number of vibrational excitations created in an optical transition increases. As such, one sees an optimal value of $S$ in this regime where processes exciting one vibrational mode become dominant. In contrast, if the vibrational mode is treated as a pure dephasing term, such a reordering phenomenon does not occur.
@@
The risetime behavior for different values of $\theta$ is discussed in Appendix B. We find that there is an initial rapid rise to a higher photon number for large $\theta$ than for a small one. That is, the photon number rise is faster for a smaller amount of electronic excitation (smaller inversion). This is a curious behavior, opposite, for example, to usual pulsed lasing, where the pulse build-up time becomes shorter for a higher amount of excitation [55]. This is because coherence plays a crucial role in SR, and larger $\theta$ (up to $\pi/2$) means higher initial coherence.
@@ IV Conclusions
Using our exact method, we benchmarked mean-field and second-order cumulant solutions, verifying their accuracy in capturing early-time dynamics when the initial state has non-zero initial coherence. This justified the application of mean-field theory to study the effect of the vibrational environment of organic emitters on the dynamics. Here we compared two models: a Holstein–Tavis–Cummings model including a discrete vibrational mode for each emitter, and a simplified phenomenological model where the vibronic coupling is replaced by Markovian pure dephasing. We found that for small vibrational coupling or dephasing, both models predict an exponential rise of the photon number at early times. In contrast, strong vibrational coupling leads to qualitatively different photon dynamics in the case of vibrational dressing, which the pure dephasing model cannot capture. Analysis of the electronic coherence revealed that sufficiently strong coupling to vibrations (roughly, $S\geq 0.3$) suppresses dynamical superradiance. As the values of $S$ in typical organic molecules are at a maximum around $0.1$, dynamical superradiance should be possible. On the other hand, the maximal values of the order $0.1$ are large enough to see specific effects of the vibrational coupling, such as the asymmetry of the risetime as a function of cavity detuning.
@@
Data availability: The datasets generated and analysed during the current study are available in the Zenodo repository at https://doi.org/10.5281/zenodo.17542025.
Appendix A Convergence of mean field approaches
Figure 7: Sum-squared error of the normalized photon number and the two-level system excitation for both mean-field and second-order cumulant solutions as defined in Eq. (13) for (a) $t_{\text{max}}=130\,\mathrm{fs}$ and (b) $t_{\text{max}}=10\,\mathrm{fs}$. For the mean-field solution $\langle\hat{n}\rangle=|\langle\hat{a}\rangle|^{2}$ and $\langle\hat{\sigma}^{+}\hat{\sigma}^{-}\rangle=|\langle\hat{\sigma}^{-}\rangle|^{2}$. The number of two-level systems ranges from $N=5$ to $N=100$ and the superradiance angle is $\theta=\pi/4$. All other parameters are the same as in Fig. 2.
In this appendix, we quantify the convergence of mean-field (mf) and second-order cumulant (c2) solutions to the exact solution for the dissipative Tavis-Cummings model from Sec. II for superradiance initial conditions ($\theta=\pi/4$). To this end, we define the sum-squared error between the mean-field or cumulant solution and the exact solution for a given quantity $\langle\hat{x}\rangle$ as
$e_{\alpha}(\hat{x})=\sum_{i=1}^{M}(\langle\hat{x}(t_{i})\rangle_{\alpha}-\langle\hat{x}(t_{i})\rangle_{\text{PIBS}})^{2},$ (13)
where $\alpha\in\{\text{mf},\text{c2}\}$. The upper bound $M$ of the sum is given by a maximum time $t_{\text{max}}=t_{M}$ up to which the error is calculated. Figure 7(a) shows the so-defined error for the normalized photon number $\langle\hat{n}\rangle/N$ and for the two-level system population $\langle\hat{\sigma}^{+}\hat{\sigma}^{-}\rangle$ for different values of $N$ ranging from $5$ to $100$. The maximum time is $t_{\text{max}}=130\,\mathrm{fs}$, which covers the entire domain calculated in Fig. 2. For increasing $N$, the mean-field error monotonically decreases for both calculated quantities. However, this is not the case for the cumulant solution, which does not exhibit a clear convergence behavior. This may at first seem surprising, but the apparent lack of convergence lies in the choice of $t_{\text{max}}$. As can be seen in Fig. 2(b) and 2(d), the cumulant solution agrees well with the exact solution up to some time $t_{0}$, after which the cumulant solution starts having irregular oscillations. It is in this region $t>t_{0}$ where a large error can build up in the cumulant solution, because the photon number oscillations can be out of phase with the exact solution. In contrast, the mean-field solution shows regular oscillations in the photon number for all times [cf. Figs. 2(b), 2(d)]. To show the convergence in the initial region $t<t_{0}$, we set $t_{\text{max}}=10\,\mathrm{fs}$ in Fig. 7(b) and calculate the same error quantities as in Fig. 7(a). The error for the photon number and for the two-level system population are equal in that case, and for both the mean-field solution and the cumulant solution a clear convergence behavior is seen for increasing $N$. These results reinforce the main message from Sec. II, namely that in the large $N$ limit, the mean-field approximation is reasonable for superradiance initial conditions.
Appendix B Influence of $\theta$ on dynamical SR